OpenStudy (anonymous):

percent of change w to 7/10 w

OpenStudy (anonymous):

30

OpenStudy (anonymous):

First divide 7/10 then multiply by 100. 7/10 = .7(x 100%) = 70%

OpenStudy (anonymous):

\[\frac{\frac{7}{10}w-w}{w}=\frac{\frac{-3}{10}w}{w}=\frac{-3}{10}=-30percent\]

OpenStudy (anonymous):

change is positive

OpenStudy (anonymous):

how so? going from a larger number to a smaller number?